OpenStudy (anonymous):

The least degree of the polynomial with real coefficients and with roots 2+i, 2i – 1 is 2

OpenStudy (wolf1728):

So what is the question?

OpenStudy (anonymous):

this is true or false

OpenStudy (wolf1728):

Yes it is true

OpenStudy (anonymous):

reason

OpenStudy (wolf1728):

Well I could try generating the original equation.

OpenStudy (wolf1728):

ROOTS = 2 + i and 2i -1 Is that right?

OpenStudy (anonymous):

kk... i had send my assignment to u would u please solve this

OpenStudy (wolf1728):

you want me to solve this right?

OpenStudy (anonymous):

plz sove this

OpenStudy (anonymous):

yes plz i need your help...

OpenStudy (wolf1728):

(x -2 -1) * (x +1 -2i) = 0 x^2 +x -2ix -2x -2 +4i -x -1 +2i =0 x^2 -2x -2ix +6i -1 = 0 Well that's what I have so far

OpenStudy (wolf1728):

Are you absolutely sure of the roots? Could you enter those again?

OpenStudy (anonymous):

next question it is true or false, give reason... If a statement has a direct proof, then it cannot be proved by contradiction.

OpenStudy (wolf1728):

I'm working on the first question

OpenStudy (mathmath333):



OpenStudy (mathmath333):

u have written -1 in place of -i

OpenStudy (wolf1728):

Yes I see that

OpenStudy (anonymous):

next question it is true or false, give reason... If a statement has a direct proof, then it cannot be proved by contradiction.

OpenStudy (wolf1728):

(x -2 -i) * (x +1 -2i) = 0 x^2 +x -2ix -2x -2 +4i -ix -i +2i^2 =0 x^2 -x -3ix -3i + 2i^2 -2 =0 Is that right?

OpenStudy (anonymous):

ya... answer next

OpenStudy (anonymous):

it is true or false, give reason... If a statement has a direct proof, then it cannot be proved by contradiction.

OpenStudy (wolf1728):

2i^2 = -2 So x^2 -x -3ix -3i -4 =0

OpenStudy (anonymous):

next question it is true or false, give reason... If a statement has a direct proof, then it cannot be proved by contradiction.

OpenStudy (xapproachesinfinity):

false my dear

OpenStudy (xapproachesinfinity):

a statement can be proven using direct proof and contradiction reasoning or even sometimes with other proves